OpenStudy (anonymous):

Green's Theorem Question! If you have a triangle that intersects these points: (0,0), (4,5), (0,5) and goes back to (0,0) again, when you are trying to find the bounds for a double integral, what do you do? I know that one of the bounds is from 0 to 5, and the other starts with zero, but I don't know how to get the upper bound.

OpenStudy (mr.math):

I think the limits are \(0<x<\frac{4}{5}y\) and \(0<y<5\).

OpenStudy (anonymous):

Hmm...I have been working on this and I think you are right, except for I think it is supposed to be 0<x<5/4y. Rise/run, yes?

OpenStudy (mr.math):

I don't think so. You can write it as \(0<x<4\) and \(0<y<\frac{5}{4}x\), it's the same thing. The line that bounds the region from the right is y=(5/4)x.

OpenStudy (anonymous):

Oh, okay, so I had y= (5/4)x, but when you have that, you have to integrate by dx first?

OpenStudy (mr.math):

It should be something like this, if you go with the second option: \[\int\limits_0^4 \int\limits_0^{\frac{5}{4}x} f(x,y)dydx\]

OpenStudy (mr.math):

Does this make sense to you?

OpenStudy (anonymous):

Yes, thanks so much for helping! I am finally doing the integral right XD

OpenStudy (mr.math):

Best of luck :)